### v3

# Repetition Threshold for Binary Automatic Sequences

Abstract
The critical exponent of an infinite word $\bf x$ is the supremum, over all finite nonempty factors $f$, of the exponent of $f$. In this note we show that for all integers $k\geq 2$, there is a binary infinite $k$-automatic sequence with critical exponent $\leq 7/3$. The same conclusion holds for Fibonacci-automatic and Tribonacci-automatic sequences.
1 Introduction
Let $w=w[1..n]$ be a finite word of length $n$. We say $w$ has period $p$, for $1\leq p\leq n$, if $w[i]=w[i+p]$ for $1\leq i\leq n-p$. We define the exponent of $w$, $\exp(w)$, to be $n/p$, where $p$ is the smallest period of $w$. Thus, for example, the French word entente has periods $3,6,7$ and exponent $7/3$.
Let $x$ be a finite or infinite word, and $\alpha>1$ be real number. If $x$ has no finite factors of exponent $\geq\alpha$, we say it is $\alpha$-free. If it has no finite factors of exponent $\alpha$, we say it is $\alpha^{+}$-free. We define the critical exponent of $\bf x$, written $\operatorname{ce}({\bf x})$, to be the supremum, over all finite nonempty factors $w$ of $\bf x$, of $\exp(w)$.
Let $S$ be a nonempty set of infinite words (or infinite sequences; we use the two terms indifferently). We define the repetition threshold $\operatorname{rt}(S)$ to be the infimum, over all elements $\bf x$ of $S$, of $\operatorname{ce}({\bf x})$. Determining the repetition threshold of a class of words is an intriguing and often difficult problem. In its most famous incarnation, we take $S$ to be the set of all infinite words over a $k$-letter alphabet $\Sigma_{k}=\{0,1,\ldots,k-1\}$. In this case we define $\operatorname{RT}(k)=\operatorname{rt}(\Sigma_{k}^{\omega})$.
Thue [25, 26] showed that $\operatorname{RT}(2)=2$. The classical example of a word achieving this bound is ${\bf t}=01101001\cdots$, the Thue-Morse sequence, which is the infinite fixed point of the morphism $0\rightarrow 01$, $1\rightarrow 10$. This word contains words of exponent $2$, but none of any higher exponent. Dejean proved that $\operatorname{RT}(3)=7/4$ and gave a very influential conjecture about the values of $\operatorname{RT}(k)$ for $k\geq 4$. This conjecture was studied by many researchers, and finally resolved positively by Currie and Rampersad [9] and Rao [22], independently.
Instead of studying the repetition threshold for all infinite words over a $k$-letter alphabet, we can also study various subsets of interest. For example, this was done for the case of Sturmian words [4], circular words [14], balanced words [21], sequences rich in palindromes [8], and episturmian words [13].
In this note we consider the repetition threshold for a very interesting class of sequences, the automatic sequences. A sequence $(a_{n})_{n\geq 0}$ is $k$-automatic if there is a deterministic finite automaton with output (DFAO) that, on input $n$ expressed in base $k$, computes $a_{n}$. In this paper, all automata read the representation of numbers starting with the most significant digit. We let $\operatorname{{\cal A}}_{k}$ denote the set of all binary $k$-automatic sequences. For more about these sequences, see [2]. For example, the Thue-Morse sequence is $2$-automatic.
We can also consider a more general notion, the generalized automatic sequences, where base $k$ is replaced by some linear numeration system. For each numeration system, one can ask for the corresponding repetition threshold for binary sequences. In general, these problems seems quite difficult. However, if the numeration system is addable—that is, the addition relation is computable by a finite automaton—then we have more tools at our disposal. Examples of this type of numeration system include the Zeckendorf system based on Fibonacci numbers, and an analogous system based on the Tribonacci numbers.
In this note we focus on the case of alphabet size $2$. Here we have a useful tool, a theorem that characterizes $\alpha$-free finite and infinite words [15]. There are three main results: Theorem 1, which exactly characterizes the repetition threshold for ${\cal A}_{k}$, $k\geq 2$; Theorem 11, which does the same thing for the binary Fibonacci-automatic sequences; and Theorem 13 for Tribonacci-automatic sequences.
2 Results for $k$-automatic sequences
For us a morphism is a map $h:\Sigma_{i}^{*}\rightarrow\Sigma_{j}^{*}$ satisfying the identity $h(xy)=h(x)h(y)$ for $x,y\in\Sigma_{i}^{*}$. If the image of each letter is of length $k$, we say that $h$ is $k$-uniform. A $1$-uniform morphism is called a coding. A morphism $h$ is called squarefree if $h(x)$ is squarefree for all squarefree $x$.
If $h(a)=ax$ for some letter $a$, and $h^{t}(x)\not=\epsilon$ for all $t$, then $h$ generates an infinite word starting with $a$ by iteration: $h^{\omega}(a)=\lim_{t\rightarrow\infty}h^{t}(a)$. Notice that a morphism can generate a squarefree word by iteration, without the morphism itself being squarefree.
Our first main result characterizes the repetition threshold for binary $k$-automatic sequences.
Theorem 1.
  • (a)
    If $k$ is a power of $2$, then $\operatorname{rt}(\operatorname{{\cal A}}_{k})=2$.
  • (b)
    If $k$ is not a power of $2$, then $\operatorname{rt}(\operatorname{{\cal A}}_{k})=7/3$.
We will need a few lemmas.
Lemma 2.
The sequence $\bf a$} is the image, under a coding, of a fixed point of a $k$-uniform morphism if and only if $\bf a$ is $k$-automatic.
Proof.
See [6, Theorem 3] or [2, Theorem 6.3.2]. ∎
Lemma 3.
Let $\bf a$} be a $k$-automatic sequence over $\Sigma_{k}$, and let $h$ be a uniform morphism from $\Sigma_{k}^{*}$ to $\Sigma_{\ell}^{*}$. Then $h({\bf a})$ is also $k$-automatic.
Proof.
See [2, Corollary 6.8.3]. ∎
Lemma 4.
Let $i,j\geq 1$ be integers and $k\geq 2$. Then a sequence is $k^{i}$-automatic if and only if it is $k^{j}$-automatic.
Proof.
See [5] or [2, Theorem 6.6.4]. ∎
Lemma 5.
Let $\bf a$} be a squarefree sequence over the alphabet $\Sigma_{4}$, and define
$h(0)$ $=011010011001001101001;$
$h(1)$ $=100101100100110010110;$
$h(2)$ $=100101100110110010110;$
$h(3)$ $=011010011011001101001.$
Then $h({\bf a})$ is $(7/3)^{+}$-free.
Proof.
See [15, Lemma 8]. ∎
Lemma 6.
There is a $k$-uniform squarefree morphism over $\Sigma_{3}$ for all $k\geq 13$, except for $k=14,15,16,20,21,22$.
Proof.
See Currie [7]. ∎
Lemma 7.
Let $\bf x$} be an infinite $7/3$-free word over $\{0,1\}$. Then ${\bf x}=p\mu({\bf y})$ for some $p\in\{\epsilon,0,00,1,11\}$ and some infinite $7/3$-free word $\bf y$}.
Proof.
This is [20, Theorem 1], which is in turn a consequence of the finitary result [15, Theorem 6]. ∎
Corollary 8.
Let $\bf x$} be an infinite $7/3$-free word over $\{0,1\}$. Then $\bf x$} contains the factors $\mu^{i}(0)$ and $\mu^{i}(1)$ for all $i\geq 0$.
Proof.
This follows by iteration of Lemma 7. ∎
Lemma 9.
The lexicographically least infinite $7/3$-free word over $\{0,1\}$ starting with $1$ is $\overline{\bf t}$ (i.e., the complement of the Thue–Morse word $\bf t$}).
Proof.
Let $\bf x$} be the lexicographically least infinite $7/3$-free word over $\{0,1\}$ starting with $1$. By Lemma 7, we have
${\bf x}$ $=1001011001101001\cdots$
$=\mu(10010110\cdots)$
$=\mu({\bf x^{\prime}}),$
where $\bf x^{\prime}$} is $7/3$-free. However, the morphism $\mu$ preserves lexicographic order, so we must have ${\bf x^{\prime}}={\bf x}$. Hence, we have ${\bf x}=\mu({\bf x})$, which implies that ${\bf x}=\overline{\bf t}$. ∎
For more on lexicographically extremal $7/3$-free words, see [20].
We recall the notion of orbit closure of an infinite sequence $\bf x$: this is the set of all infinite sequences $\bf y$ such that every finite prefix of $\bf y$ appears somewhere in $\bf x$.
Lemma 10.
Let $\bf x$} be a generalized automatic sequence in an addable numeration system and let $w$ be a finite word. The lexicographically least sequence in the orbit closure of $\bf x$} beginning with $w$ is also a generalized automatic sequence for the same numeration system as $\bf x$}.
Proof.
This can be shown using Walnut (a theorem-prover for automatic sequences [17, 24]) by an easy modification of the approach described in [1] or [24, Theorem 8.11.1]. ∎
We can now prove our first main result, Theorem 1.
Proof.
(a) If $k$ is a power of $2$, then by Lemma 4, a $2$-automatic sequence is also a $2^{i}$-automatic sequence for all $i\geq 1$. So it suffices to prove the result for $k=2$. But then $\bf t$ is a $2$-automatic sequence that avoids overlaps, and it is easy to see that no binary word of length >3>3 can avoid squares.
(b) Suppose $k$ is not a power of $2$. Suppose there is some ${\bf x}\in\operatorname{{\cal A}}_{k}$ with $\operatorname{ce}({\bf x})=\alpha<7/3$. Then, by Corollary 8, we know that $\bf x$ contains $\mu^{i}(0)$ as a factor for arbitrarily large $i$. Therefore the orbit closure of $\bf x$ contains all sequences in the orbit closure of $\bf t$. By Lemma 9, the lexicographically smallest $7/3$-free sequence starting with $1$ is $\overline{\bf t}$, and hence the lexicographically smallest sequence in the orbit closure of $\bf x$} starting with $1$ is also $\overline{\bf t}$. It now follows from Lemma 10 that $\overline{\bf t}$ is $k$-automatic. Then, by a theorem of Cobham [5], we know (since $k$ and $2$ are multiplicatively independent) that $\overline{\bf t}$ must be ultimately periodic, a contradiction.
Therefore $\operatorname{ce}({\bf x})\geq 7/3$. It remains to show that there is a word of critical exponent $7/3$ for all the relevant $k$.
By combining Lemmas 6, 5, and 3, we see that the desired result holds for all $k\geq 23$, and also for $k=13,17,18,19$. Thus, the only remaining cases are
$k=3,5,6,7,9,10,11,12,14,15,20,21,22.$
We now treat each of these cases. The strategy is to find, for each of these values of $k$, a $k$-uniform morphism over $\Sigma_{4}$ whose fixed point is squarefree, and then apply Lemma 5. In some cases, the morphism we find is actually over $\Sigma_{3}$. It is not necessary to treat $k=9$, since it is implied by the result for $k=3$ by Lemma 4.
We found the appropriate morphisms through either breadth-first or depth-first search. Table 1 gives the images of the morphisms we found.
$k$ morphism
3 $0\rightarrow 021$
$1\rightarrow 031$
$2\rightarrow 012$
$3\rightarrow 013$
5 $0\rightarrow 01321$
$1\rightarrow 01231$
$2\rightarrow 02031$
$3\rightarrow 02321$
6 $0\rightarrow 013121$
$1\rightarrow 013031$
$2\rightarrow 013231$
$3\rightarrow 012131$
7 $0\rightarrow 0123021$
$1\rightarrow 0123121$
$2\rightarrow 0120321$
$3\rightarrow 0121321$
10 $0\rightarrow 0120131021$
$1\rightarrow 0120231021$
$2\rightarrow 0120301021$
$3\rightarrow 0120312021$
11 $0\rightarrow 01201020121$
$1\rightarrow 01202120121$
$2\rightarrow 02010212021$
12 $0\rightarrow 012021201021$
$1\rightarrow 012010212021$
$2\rightarrow 012102120121$
14 $0\rightarrow 01201023201021$
$1\rightarrow 01201030201021$
$2\rightarrow 01201031201021$
$3\rightarrow 01201020301021$
15 $0\rightarrow 012010230201021$
$1\rightarrow 012010231201021$
$2\rightarrow 012010203201021$
$3\rightarrow 012010213201021$
20 $0\rightarrow 01201020120210201021$
$1\rightarrow 01201021201210201021$
$2\rightarrow 01201020121021201021$
21 $0\rightarrow 012010201230210201021$
$1\rightarrow 012010201231210201021$
$2\rightarrow 012010201203210201021$
$3\rightarrow 012010201213210201021$
22 $0\rightarrow 0120102012030210201021$
$1\rightarrow 0120102012130210201021$
$2\rightarrow 0120102012320210201021$
$3\rightarrow 0120102012301210201021$
Table 1: Morphisms generating squarefree words by iteration.
We can prove that each of these words are squarefree using the following model for $k=3$:
morphism ff3 "0->021 1->031 2->012 3->013":
promote FF3 ff3:
eval test3 "?msd_3 ~Ei,n (n>0) & Au,v (u>=i & u<i+n & v=u+n) => FF3[u]=FF3[v]":
And Walnut returns TRUE for all of them. ∎
Since the construction involving the morphism $h$ does not necessarily produce an automatic sequence with the smallest possible DFAO for a $k$-automatic $(7/3)^{+}$-free infinite binary word, it may be of interest to try to determine the number of states in a smallest DFAO. Lower bounds can be found by breadth-first search; upper bounds can be found by applying the construction involving $h$ or other means. Table 2 summarizes what we currently know. An asterisk denotes those entries where the exact value is known; intervals denote known lower and upper bounds. The automata can be constructed straightforwardly from the expression as the iteration of a fixed point starting with 0, followed by the specified coding.
number of
$k$ states $s$ automaton
22 22* $0\rightarrow 01$
$1\rightarrow 10$
33 $8\leq s\leq 16$ See below.
44 22* $0\rightarrow 0110$
$1\rightarrow 1001$
55 44* $0\rightarrow 01230$ (0)
$1\rightarrow 20120$ (0)
$2\rightarrow 23012$ (1)
$3\rightarrow 02302$ (1)
66 $6\leq s\leq 27$ Apply $h$ to word generated by
morphism for $k=6$ in Table 1.
77 44* $0\rightarrow 0120330$ (0)
$1\rightarrow 1211321$ (0)
$2\rightarrow 2112033$ (1)
$3\rightarrow 3013303$ (1)
88 22* $0\rightarrow 01101001$
$1\rightarrow 10010110$
99 44* $0\rightarrow 012301312$ (0)
$1\rightarrow 312012312$ (0)
$2\rightarrow 301312312$ (1)
$3\rightarrow 013123012$ (1)
$10$ 44* $0\rightarrow 0120132132$ (0)
$1\rightarrow 0121321201$ (0)
$2\rightarrow 3212012132$ (1)
$3\rightarrow 3201213201$ (1)
$11$ 44* $0\rightarrow 01202301230$ (0)
$1\rightarrow 20120230201$ (0)
$2\rightarrow 23012012302$ (1)
$3\rightarrow 30123020123$ (1)
Table 2: Minimum number of states needed to produce a $k$-automatic $(7/3)^{+}$-free infinite binary word.
To get a $16$-state DFAO D3 for base $3$, use the following Walnut commands:
morphism d9 "0->012301312 1->312012312 2->301312312 3->013123012":
promote D d9:
morphism code "0->0 1->0 2->1 3->1":
image D9 code D:
eval test9a "?msd_9 ~Ei,p (p>0) & Au (u>=i & 3*u<=3*i+4*p) => D9[u]=D9[u+p]":
# tests that morphism does indeed generate a (7/3)+-free word
convert D3 msd_3 D9:
These commands take a DFAO for a $9$-automatic word and compute a DFAO for inputs in base $3$.
3 Fibonacci- and Tribonacci-automatic sequences
In the Zeckendorf or Fibonacci numeration system [16, 27], we write every natural number $n$ uniquely as a sum of Fibonacci numbers $F_{i}$ for $i\geq 2$, subject to the condition that we never use both $F_{i}$ and $F_{i+1}$ in the sum. The representation can be viewed as a binary word, where we write $1$ if $F_{i}$ is included and 0 otherwise. A sequence $(a_{n})_{n\geq 0}$ is said to be Fibonacci-automatic if there is a DFAO that on input $n$ expressed in the Zeckendorf numeration system reaches a state with output $a_{n}$ [18, 10, 11]. We let $\operatorname{{\cal A}}_{F}$ denote the set of all binary Fibonacci-automatic sequences. The following is our second main result.
Theorem 11.
We have $\operatorname{rt}(\operatorname{{\cal A}}_{F})=7/3$.
Proof.
Let $\bf x$ be Fibonacci-automatic. The possibility $\operatorname{ce}({\bf x})<7/3$ is ruled out by the same argument as in the proof of Lemma 10 (b) above. Namely, the automaton gives $\bf x$ as the image, under a coding, of a fixed point of some morphism $\varphi$ (as in Example 1 in [23, Section 5]). As in the base-$k$ case, the dominant eigenvalue of the incidence matrix of $\varphi$ is $\alpha=\frac{1}{2}(1+\sqrt{5})$. An easy induction proves that $\alpha^{i}=L_{i}/2+F_{i}\sqrt{5}/2$, where $L_{i}$ is the $i$’th Lucas number, so $\alpha$ and $2$ are multiplicatively independent. Then a theorem of Durand [12] finishes this part of the proof.
So it suffices to give an example achieving the exponent $7/3$. The easiest way to do this is to find a Fibonacci-automatic squarefree sequence $\bf y$ over a $4$-letter alphabet, and then apply the morphism $h$ to $\bf y$. Such a sequence is computed by the DFAO in Figure 1.
Figure 1: Fibonacci automaton computing the sequence $\bf y$.
Once we have saved the DFAO in Figure 1 in the Word Automata directory of Walnut under the name Y.txt, we can verify that the sequence it generates is squarefree, as follows:
eval ysquf "?msd_fib ~Ei,n n>=1 & At (t<n) => Y[i+t]=Y[i+t+n]":
and Walnut returns TRUE. ∎
Remark 12.
By the way, the critical exponent of the word y is $(5+\sqrt{5})/4\doteq 1.809$, although we do not use this fact anywhere. The reader can compare y to the word ${\bf x}_{4}$ in [21].
When we apply the morphism $h$ to $\bf y$, we get a DFAO of 943 states. It can be computed via the following Walnut code:
morphism ks
 "0->011010011001001101001 1->100101100100110010110
  2->100101100110110010110 3->011010011011001101001":
image YP ks Y:
This naturally suggests the question, is there a smaller Fibonacci DFAO computing a $(7/3)^{+}$-free binary sequence?
There is such a sequence ${\bf w}=01101001101100110100\cdots$ generated by a DFAO of $32$ states. Let $x^{R}$ denote the reversal of the finite word $x$. We can construct $\bf w$ iteratively as follows: define
$w_{0}$ $=011$
$w_{1}$ $=01101$
$w_{n}$ $=\begin{cases}w_{n-1}\overline{w_{n-2}}^{R},&\text{if $n\geq 2$ and $n\equiv 1,2\ \mbox{\rm(mod}\ 3\mbox{\rm)}$};\\[7.22743pt] w_{n-1}w_{n-2}^{R},&\text{if $n\geq 0\ \mbox{\rm(mod}\ 3\mbox{\rm)}$}.\end{cases}$
Clearly each $w_{i}$ is a prefix of all $w_{j}$ for $j\geq i$, so we can define $\bf w$ as the unique infinite word of which all the $w_{i}$ are prefixes. It is not difficult to show, with Walnut, that this word is generated by a Fibonacci DFAO with $32$ states and is $(7/3)^{+}$-free. We omit the details. This means the size of the smallest such DFA is at most $32$. Breadth-first search shows it is at least $12$.
One can also carry out the same sorts of computations for the Tribonacci-automatic sequences, where we use a numeration system based on the Tribonacci numbers [19]. This gives our final main result.
Theorem 13.
The repetition threshold for binary Tribonacci-automatic sequences is $7/3$.
Proof.
The proof is quite similar to the proof of Theorem 11, so we just sketch the details.
By breadth-first search, we can identify a candidate squarefree Tribonacci-automatic sequence $\bf z$ over $\Sigma_{4}$, and prove it is indeed squarefree using Walnut. It is generated by the DFAO in Figure 2.
Figure 2: Tribonacci automaton computing a squarefree sequence.
We can then apply the morphism $h$ to obtain a $(7/3)^{+}$-free binary Tribonacci-automatic sequence generated by a DFAO with 36849 states. No binary sequence of smaller exponent is possible, by an argument similar to that in the proof of Theorem 11.
Remark 14.
It can be shown that the critical exponent of the sequence $\bf z$ is $1.875954364915\cdots$, the real zero of the polynomial $4X^{3}-24X^{2}+56X-47$.
4 Going further
It seems likely that $7/3$ is the critical exponent for many other binary generalized automatic sequences, such as the Pell-automatic sequences [3], but a general theorem seems hard to prove.
Acknowledgments
We gratefully acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). Nous remercions le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG) de son soutien. The research of NR was supported by NSERC, grants RGPIN-2019-04111 and RGPIN-2025-04076. The research of JS was supported by NSERC, grant RGPIN-2024-03725.